find the total surface area of a rectangular solid with a lenght of 9 ft a width of 11 ft and a height of 14 ft.

Found 2 solutions by kietra, sarah_adam:
Answer by kietra(57)

To find the surface area of a 3D object, you need to basically find the area of each of the faces of the figure. Since the solid is rectangular, some of the sides will have the same surface area. The top and bottom, front and back, and then both sides.
The area of top and bottom (depending on how you label sides) would be 2(11x14)=2(154)=308.
The area of the front and back would be 2(9x11)=2(99)=198.
And the area of the sides would be 2(14x9)=2(126)=252.
Add all the sides together 308 + 198 + 252 = 488.

Answer by sarah_adam(201)

Given l =9ft , w = 11ft and h = 14 ft
Total surface area = 2{(lw)+(wh)+(lh)}
2{(9*11)+(11*14)+(9*14)}= 958 sq ft